OpenStudy (yacoub1993):

How many prime numbers are there between 21 and 35? A. 3 B. 4 C. none D. 2

OpenStudy (mathstudent55):

23, 29, 31

OpenStudy (anonymous):

A. 3

OpenStudy (anonymous):

@genius12 >Implying that Yacoub1993 can't count.

OpenStudy (anonymous):

prime= numbers that can only be multiplied by 1 and itself such as (23)(1) = 23 (29)(1) = 29 (31)(1) = 31 the rest are composite numbers :)

OpenStudy (yacoub1993):

thnxs guys I already figured it out :)))

OpenStudy (mathstudent55):

@some_someone prime= numbers that can only be multiplied by 1 and itself Not quite. prime= number that can only be _divided evenly_ by 1 and itself

OpenStudy (anonymous):

yeah, but its sort of the same concept, so its not really wrong either lol :) but yeah urs is more precise :D

OpenStudy (mathstudent55):

@some_someone Every number can be multiplied by 1 and itself, so according to you, every number is prime. I don't think that is the definition of a prime number.

OpenStudy (anonymous):

noooo, i said ****ONLY**** not every!! please get ur facts straight before accusing me thanx :)

OpenStudy (mathstudent55):

Ok, then name a number that can only be multiplied by 1 and itself and then name a number that cannot.

OpenStudy (anonymous):

ok yeah 23 * 1 = 23 only could multiply by 1 and itself (23) :) 27 is composite since it can be multiplied by 27 *1 = 27 and 9 * 3 = 27

OpenStudy (mathstudent55):

Ok, 23 * 1 = 23 23 * 2 = 46 23 * 3 = 69 23 * 4 = 92 23 * 5 = 115 etc. I thought you said 23 can only be multiplied by 1 and itself, but I just multiplied it by all whole numbers from 1 to 5.

OpenStudy (anonymous):

yes but it could only be multiplied by one to get 23! you got just multiply like 11 * 2 = 23; its not right.